Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
gouaty, letendre, masserot - 16:21 Friday 09 April 2021 (51372) Print this report
Comment to DRMI recovery: lock with 1f signals & PR alignment in full bandwidth, possible issue with 3f signals (51361)
At 14h02m18-UTC we reconfigured SIB2_DEMOD_B2_PD1_2 in SIB2_dbox_bench.
This solved the problem. The B2 PD1 RF channel is now behaving properly.
Search Help